OpenStudy (anonymous):

At a town fair, 30 tickets cost $20 How many tickets could you expect to receive for $8, assuming there is no discount for buying more tickets?

OpenStudy (tanner23456):

set up a proportion: 30/20 = x/8 cross multiply 20x = 240 now just solve for x